Most

Version 1.56 23 April 1994

Most is a 2.0 only textviewer which I wrote with full workbench support in mind. It supports ARexx, help via AmigaGuide, locale.library, XPK, AppWindows, AppIcons, Clipboard, Pipes, no limited number of windows and more. Most is GiftWare.

by Uwe Röhm

Documentation for Most V1.56 Date: 23 April 1994 Author: Uwe Röhm

Most is a 2.0 only textviewer which I wrote with full workbench support in mind. First at all an overview of interesting features:

− unlimited number of windows − each text has its own process − search function − highly configurable − clipboard support − “click-and-load” function for includes − font sensitive GUI − PubScreen, CustomScreen or Workbench − AppWindow, AppIcon − ARexx Port − localization with locale.library − AmigaGuide help files − XPK support − pipe support − as mentioned, kickstart 2.0 is requiered

1 Copyright

Most 1.56 is published under the concept of GiftWare! So if you find Most a useful tool (and of course use it) please make me a present of any kind - and make me very happy.

Beside this I keep my Copyright:

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1992/93/94 by Uwe Röhm

You must get my aggreement if you wish to publish Most in a commercial product. Especially it is not allowed to use Most for military purposes of any kind which includes using Most in companies which plan or only support military products.

You are allowed to include Most in PD packages which cost less than $10.- ! But I ask you to keep the archive together. You have to include the documentation with each version. Thank you!

2 Background

There are some goals I try to achieve with Most:

− full workbench support − easy to use with mouse and keyboard − flexibility − memory friendly :-)

So Most uses a window for each text - there is no possibility to use it like Commodore’s More in a shell. Of course you can choose another screen to open on or even an own CustomScreen. How much windows are opened by Most and where they are is absolut without interest and only limited by the amount of free (chip-) memory!

As each text is displayed in his own window there is also its own process behind. Cause of the windows are absolut independent of each other. Especially you can send ARexx commands to different windows simultaneous and you are free to open new Most windows even if previously opened windows are “busy” - there is Most’s main process which handles all texts and some global ARexx commands (see Chapter 11 [ARexx Port], page 22).

To be memory friendly and to display very huge texts even on a 500KB system Most does buffering. Small texts, which fit completly into the buffer, are read only once and closed than so that you can e.g. edit and save it again. But if the text is larger than the configurated buffer length (yes, of course you can change the buffer size, but read on :-) ) Most will do some sort of “paging”. So memory usage will be limited by the buffer size you specified (little memory -> small buffer). On the other hand this slowes scrolling and searching down!!! Sorry for this, but if you have enough memory just increase to buffer size to - let’s say - 2 MB. Don’t be afraid. Of course Most only uses the amount of memory it needs if a text is smaller than the buffer ;-)

There is one exception of the paging method: XPK packed files. In the current version Most does no paging to XPK files but reads the complete file into memory (see Chapter 9 [XPK support], page 20). Perhaps this will change in future, perhaps it will stay…

Oh, I nearly forgot that Most can display all ANSI sequences. Without this it wouldn’t be a real text viewer. As I don’t want to do old work again and again I decided to use the console.device. This enables you to mark lines and copy them with Amiga C into the clipboard! I regret that there are also some problems:

− You can only mark text inside the visible text portion. − Scrolling is quiet annoying on an Amiga without 680x0 (x > 1). − Don’t load binary files. I warned you!!!

3 Configuration

Most does the following configuration steps on startup:

1. internal defaults

2. Environments Most/Most and Most/FileReq (When used from shell local environments will be read first!)

3.

Shell-start: parse command line WB-start : parse ToolTypes

4. NEW since Most 1.53: For each text loaded Most will read the project icon (if exists) and use the new tooltypes. You can also include all ARexx commands in the tooltypes which will be executed just after loading the new text! So you can write local ARexx scripts directly into the icons of texts. For a list of the possible commands see Chapter 11 [ARexx Port], page 22.

And now, ladys and gentlemen, here are the options to be used in 2.), 3.) and 4.):

3.1 FILES

Usage A list of filenames to be displayed (only on shell command line!). You can use any AmigaDOS wildcards. Since version 1.55 Most also finds files with names, which include wildcards itself (like “(test)”).

3.2 PORTNAME

Usage Custom name of global Chapter 11 [ARexx Port], page 22 of Most. Overrides the default name.

Default MOSTPORT

3.3 PUBSCREEN

Usage Name of PubScreen where Most will open its windows on.

3.4 STARTUP

Usage Name of an ARexx script which will be executed every time Most opens a window. (see Chapter 11 [ARexx Port], page 22). Remember that you can use directly all ARexx commands as tooltypes in project icons of text files (as a sort of "tooltype script").

3.5 WINDOWX

Shortcut WX Usage X-coordinate of Chapter 6 [The main window], page 17

3.6 WINDOWY

Shortcut WY Usage Y-coordinate of Chapter 6 [The main window], page 17. -1 for centered windows

3.7 WINDOWW

Shortcut WW Usage Width of windows. If you specify negative values or 0, the width of the screen will be

used instead.

Default 600

3.8 WINDOWH

Shortcut WH Usage Height of windows. If you specify negative values or 0, the height of the screen (without

its title bar) will be used instead.

Default 400

3.9 SCREENMODE

Shortcut SM

Usage A valid screenmode Most will use to open its on screen with (which will be a PubScreen named Most). You can get a list of the avaible screenmodes of your system if you start

          Prefs/ScreenMode.

Attention! The names of the screenmodes are different for most versions of the monitor files. Remember that you must specify a valid name, which means, that you must spell it exactly as in the monitor files. If in doubt start Prefs/ScreenMode and read the listview.

Example

                  SCREENMODE "Pal:HighRes"

3.10 FONT

Usage name of a font to be used. The name must include the .font postfix. The height of the font can be adjusted with the Section 3.11 [FONTHEIGHT], page 7 option. Be careful with proportional fonts.

Example

                  FONT "helvetica.font"

3.11 FONTHEIGHT

Usage Tells Most which font height it should use. Works only together with the Section 3.10 [FONT], page 7 option. If not specified Most uses the system default font’s height.

3.12 APPICON

Usage A switch, which lets Most use an AppIcon on workbench. If you start Most with the APPICON option and without further file names, Most will only put its AppIcon onto workbench and stay in background. (see Chapter 8 [The AppIcon], page 19)

3.13 ICONX

Usage X-coordinate of Most’s Section 3.12 [APPICON], page 7 on workbench

3.14 ICONY

Usage Y-coordinate of Most’s Section 3.12 [APPICON], page 7 on workbench

3.15 BUFLEN

Usage Very interesting option to justify Most’s buffer size. In general: On small systems use small values, on larger systems try higher buffer sizes. BUFLEN defaults to 50 KB (which is quiet small :-) ). Texts smaller than the buffer size will be read complete into memory and use only the necessary amount of memory.

Default 50 000 Bytes

3.16 PIPELEN

Usage Most can also read out of a pipe. Normally it would allocate a buffer of Section 3.15 [BUFLEN], page 8/10 Bytes for reading. With this option you can justify the buffer length for pipes by hand.

Example

                  PIPELEN 8192

3.17 QUOTES

Usage You can specify a sequence of quote characters. These characters will be used for the “click-and-load” option: If you doubleclick on a word which stands between a quote character Most will try to load a text named like this word. This is very useful for programmers as you can view include files very easy!

Default "’<>(){}[]

3.18 DIRS

Usage Specify one or more directories Most will search through if you use the “click-and-load” function (see Section 3.17 [QUOTES], page 8). All directories must be seperated by |.

Example

                  DIRS "Include:|TexInputs:|LC:Include/"

3.19 ENDS

Usage Postfixes to be appended on include files with the “click-and-load” function in the given order (see Section 3.18 [DIRS], page 9, see Section 3.17 [QUOTES], page 8).

Example

                  ENDS ".c.h.tex"

3.20 STACKSIZE

Usage Adjust the size of the editor’s stack. Useful for example for the emacs ("Eight Megabytes And Constantly Swapping"?)…

Example

                  STACKSIZE 50000

Default 8192 bytes

4 Keyboard commands

In the current version Most supports all (and more!) keyboard shortcuts well known from More. Future versions will allow custom keyboard commands…

4.1 Scrolling

one line forwards

          CURSOR DOWN NUMERICPAD 2 RETURN ENTER

one line backwards

          CURSOR UP NUMERICPAD 8 DELETE

one page forwards

          SHIFT CURSOR DOWN SPACE PAGE DOWN

one page backwards

          SHIFT CURSOR UP BACKSPACE PAGE UP b

Jump to beginning of text

          HOME <

Jump to end of text

          END >

oen column to the right

          CURSOR RIGHT NUMERICPAD 6

one column to the left

          CURSOR LEFT NUMERICPAD 4

8 columns to the right

          SHIFT CURSOR RIGHT

8 columns to the left

          SHIFT CURSOR LEFT

jump to first column

          NUMERICPAD 5

4.2 Searching

search actual searchstring again

          n

ignore upper/lower case

          .

searching with distinction of upper/lower case

          /

Open the search requester

          f s

4.3 Open file

Load new text with l (a filerequester shows up).

4.4 Info requester

Beside the usual Chapter 1 [Copyright], page 2 you find here some information about the current text: Its filename, the filelength, the number of lines and the visual number of lines, and also the name of the XPK packer (or just “—-” if the file isn’t packed). At last there’s a message indicating that the current text has been read completly into memory or – if not – the length of the used read buffer.

The Info requester can be opened with h.

4.5 Help

If you press HELP, the AmigaGuide will be started with the documentation of Most. Of course this requires the AmigaGuide be installed correctly. If Most can’t open the amigaguide.library it will open the Section 4.4 [Info requester], page 13.

AmigaGuide searches its documents in the current directory and also in the paths listed in the environment variable AmigaGuide/Path.

4.6 Edit file

If you press e, the current text will be loaded into an editor. The editor will be started in the background. The name of the editor must be configurated in the environment variable EDITOR. You can adjust the editor’s stack size with the Section 3.20 [STACKSIZE], page 11 option.

4.7 Quit file

Close the text window with ESCAPE, CTRL-C or q.

5 Menus

There are the following menus available:

5.1 Project

New… Opens a new window and brings up a filerequester where you can choose a text file to be displayed in the new window. If you cancel the filerequester the window will be closed. The position and dimensions of the filerequester can be stored in Section 5.5 [Settings], page 16.

Open… Asks you for a new file to be loaded into the current window.

Edit… Starts the editor with the current text via RUN. The name of the editor can be configurated in ENV:Editor (see Section 11.11 [Edit], page 25), its stack size can be adjusted with the Section 3.20 [STACKSIZE], page 11 option.

Print Simply prints the current text to PRT:.

Info… Pops up the Section 4.4 [Info requester], page 13 where you find some useful information to the current text like name, length, line count, visible lines, XPK packername and a short comment if the text has been loaded completely into memory or if it uses an buffer.

Quit Quits the current window (and if it was the last one Most at all)

5.2 Move

Move through the current text as the menus describe.

With the menu “Jump to line...” you can specify a certain line number in a requester where you will jump to.

5.3 Menu Search

searching. . .

Chapter 7 [The search requester], page 18 will be opened with its stringgadget activated. If the search requester is open this menu is disabled.

forwards/backwards

Search string forwards/backwards (only with open search requester)

5.4 Extras

Refresh page

Reset console and redraw page

Set tabulator width. . .

Enter new numeric value for the tabulator width (per default every 8th character there’s a tab).

Execute ARexx Macro. . .

Asks for an ARexx Macro to be executed (see Chapter 11 [ARexx Port], page 22).

5.5 Settings

save
global Prefs

Save actual settings in two environments in ENV: and ENVARC:. A subdirectory Most will be created automatically. All known options from workbench or shell start are saved and also the dimensions and position and the pattern string of the file requester.

local Prefs

Does the same as above, but uses local shell variables. This only works if you started Most from a Shell!

6 The main window

Title In the title the filename and on the right hand side the current column and line numbers

are displayed. Format:

      <filename>                                  <column>/<line>

Gadgets Beside the system gadgets there are three further ones which control text scrolling: You can use the scroller continously by click-and-draging or for scrolling one page forwards/backwards if you click under/above the knob. The knob’s position represents the current text position, its height the relation between the visible and the total text length. Use the arrows to scroll one line in the correspondending direction.

Mouse “Click-And-Load” function.

If you doubleclick into the window Most will look at this position to find a word which is limited by one of the configurated quote characters (see Section 3.17 [QUOTES], page 8). This word (can include spaces!!) will be treated as a filename for which Most searches first in the current directory and than in each further configurated (see Section 3.18 [DIRS], page 9) directory. Most will also try the plain filename and than (if not found) append each prefix given with the Section 3.19 [ENDS], page 10 option. The first found file will be displayed in a new window. If no name matches, the screen will flash (DisplayBeep). Example:

Doubleclick on “#include <exec/exec.h>” between “<” and “>”.

the file Include:Exec/Exec.h will be loaded, if you

specifed “DIRS Include:” on startup.

Clipboard You can easy mark any text in the visible text by click-and-dragging (drag without releasing the mouse button). Copy the marked block with Amiga-C into the clipboard. The current version doesn’t support paste from the clipboard.

AppWindow Take any “real” icon (no AppIcon) on workbench and drop it in Most’s window. It will be loaded into the current window. If you drop more than one icons Most will open further windows.

7 The search requester

Here you can enter a search string and also manipulate the three search option. Therefor you can use the mouse and/or the keyboard shortcuts indicated by the underlined characters. Additionaly you have the same functionality as in the main window! So no annoying re-activation of Chapter 6 [The main window], page 17.

− All menus are avaible as in the text window. Only the Section 5.3 [Menu Search], page 15 will

be disabled and instead “search forwards” and “search backwards” are unghosted. − Beside the mentioned Requester shortcuts you can use all normal Chapter 4 [Keyboard com-

mands], page 12 like scrolling with the cursor keys!!! (Be sure the string gadget isn’t activated)

8 The AppIcon

The AppIcon has two functions

Doubleclick

A filerequester will come up where you can choose a file to be loaded and displayed. If you select “Cancel” Most will show a security requester where you can quit Most.

Icons Drag-and-drop one or more icons on the AppIcon and they will be loaded.

Comments:

Most uses his own program icon as its Section 3.12 [APPICON], page 7. Perhaps I will support custom icons in the future but not yet ;-) The coordinates of the AppIcon can be saved in the menu Section 5.5 [Settings], page 16. Use the options Section 3.13 [ICONX], page 8 and Section 3.14 [ICONY], page 8 for exact placement of the AppIcon.

9 XPK support

Main goal of Most’s XPK support is to hide any consequences of packed data from the user. This means, that any packed file (with XPK) will automatically be unpacked from Most if the XPK packages is correct installed. The only hint you get is in the information requester, where the line “XPK-packer:” will show you the used packer (usually you see “—-” there). There is only one thing special with XPK packed data: The file will be loaded completely into memory and no buffering will be done!

If Most finds an encrypted files (with XPK of course), it opens a password requester and prompts you for the password. Be careful, because this is your only chance! If you give the wrong answer, you will see the encrypted file and have to reload it.

On the other hand if you don’t have XPK you can use Most without any difference. The xpkmaster.library is not required by Most to run at all. Of course in this case XPK packed data will not be unpacked :-)

10 Pipe support

As a special feature Most 1.42 supports pipes. This means, that you can type something like “type foo | Most” at your (W)Shell prompt and Most will show the output of the first type command in its window!

As I said this feature is brand new. So I regret that I couldn’t test it for all possibilities. Please report me any failure.

There are two small problems:

1. XPK doesn’t work with pipes.

2. Second there is the problem with pipes that you cannot know how much data therein will be… So one effect is, that you cannot scroll back to the start of an file if you scrolled over the bounds of a buffer and the other one is that the size of the buffer will be (maximum Section 3.15 [BUFLEN], page 8 / 10) every time (you can set the pipe’s buffer size with the Section 3.16 [PIPELEN], page 8 option). If the pipe’s data fits completly into the buffer you have no restrictions for scrolling. But if the data is to large - Sorry.

11 ARexx Port

Beside the global ARexx port MOSTPORT (or the name you gave it with the option Section 3.2 [PORTNAME], page 5 on startup) of Most’s main process each text window has his own private ARexx port. To select a window first send the command Section 11.1 [GetFileList], page 22 to the global port. You will receive a list of the avaible windows and its portnames.

All following ARexx commands can be send to a private window port, but Most’s main port does only understand the first 6 commands Section 11.1 [GetFileList], page 22, Section 11.2 [GetScreenName], page 23, Section 11.3 [Quit], page 23, Section 11.6 [OpenNew], page 24, Section 11.5 [Set], page 24 and Section 11.4 [Request], page 23.

In RC Most will return an errorcode of

• 0 for no errors

• 5 if it couln’d open a new text with Section 11.7 [Open], page 24 or Section 11.6 [OpenNew], page 24

• 10 if the parameters for a command were wrong

• 20 for an unknown ARexx command

All other commands will return an errorcode of 20 if send to MOSTPORT. Sorry, but the ARexx port isn’t complete today. But I’m working on it :-)

The following ARexx commands are supported:

11.1 GetFileList

Template GetFileList

Result In RESULT you get a list of all available text windows and their port names. RESULT looks like this: {<window’s name> <port name> }^+

Example

                  GetFileList
                  RESULT: "Most.dok MOST.0 Most.doc MOST.1"

11.2 GetScreenName

Template GetScreenName Result RESULT contains screen name. Usage Get the name of the screen, where the current window is on. Example

                  GetScreenName
                  RESULT - "Workbench"

11.3 Quit

Template Quit Usage Quits Most and closes all windows. Chapter 8 [The AppIcon], page 19 of Most will be

deleted – if there is one at all.

Example

                  Quit

11.4 Request

Template Request TITLE/K/A TEXT/K/A GADGETS/K/A

Resultat RESULT contains number of selected gadget:

0 the most right gadget (should be “Cancel”)

1 bis n-1 the other gadgets from left to right

Usage A simple method of asking the user silly questions via requester. Title and text are simple strings. With GADGETS you specify the texts of the gadgets (and how much gadgets there are), seperated with ’|’.

Example

                  Request TITLE "Test" TEXT "All OK?" GADGETS "YES|No"
                  RESULT - 1 if YES has been selected, else 0

11.5 Set

Template Set CASE/S WORD/S BACK/S

Usage Set some internal parameters like CASE ignore upper/lower case on/off WORD search only for whole words on/off BACK search backwards on/off All parameters will be set if specified, and disabled otherwise.

Example

                  Set CASE ON
                  Enables casesensitiv searching, but disables the other two options.

11.6 OpenNew

Template OpenNew FILE/K Usage Load a new text in a new window. The filename is optional. If you don’t specify a

name a filerequester opens.

You don’t need to select a current window before. This command doesn’t return until the new window is closed again!

Example

                  OpenNew FILE Most.dok

11.7 Open

Template Open FILE/K Usage Load a new text in the current window. The filename is optional. If you don’t specify

a name a filerequester opens. This command doesn’t return until the window is closed again!

Example

                  Open FILE Most.dok

11.8 Close

Template Close Usage Closes the current window.

11.9 ToFront

Template ToFront Usage Makes the current visible: It will be made the frontmost window and the screen it’s on

will be activated and go to front.

11.10 GetWindowDims

Template GetWindowDims

Result  In RESULT stands <xpos> <ypos> <breite> <höhe>

Usage Get the position and dimensions of the current window in RESULT Example

                  GetWindowDims
                  RESULT - "40 40 600 180"

11.11 Edit

Template Edit Usage Section 11.11 [Edit], page 25 the current text (runs editor).

11.12 Info

Template Info Usage Opens the Section 4.4 [Info requester], page 13 to the current window.

11.13 FirstPage

Template FirstPage Usage Show the first page of the current text.

11.14 LastPage

Template LastPage Usage Show the last page of the current text.

11.15 NextPage

Template NextPage Usage Show the next page of the current text.

11.16 PrevPage

Template PrevPage Usage Show the previous page of the current text.

11.17 NextLine

Template NextLine Usage Scroll up one line.

11.18 PrevLine

Template PrevLine Usage Scroll down one line.

11.19 NextColumn

Template NextColumn Usage Scroll left one column.

11.20 PrevColumn

Template PrevColumn Usage Scroll right one column.

11.21 NextTab

Template NextTab Usage Scroll left 8 columns.

11.22 PrevTab

Template PrevTab Usage Scroll right 8 columns.

11.23 Goto

Template Goto LINE/K/N/A COLUMN/K/N

Usage Jump to the specified line. Optional you can specify a new first column. Example

                  Goto LINE 142

11.24 Search

Template Search TEXT/K Usage Without argument: Chapter 7 [The search requester], page 18 opens or if allready open

the current string will be search again.

With searchstring: Search for the given string.

Example

                  Search TEXT foobar

12 Known bugs

− Several control sequences for cursor movements trash the display. − Most has problems with the pipe of WShell. − An Enforcer-hit with Most 1.53 has been reported which may be still there. The input buffer

seems to overflow after havy input… − Bruno Lacombe reported a Mungwall hit with Most 1.53 I couldn’t reproduce here…

13 Ideas - Bug reports

As mentioned before Most isn’t complete today. There are some points I will improve in the future but I can’t promise it. Please remember that I write Most in my spare time without gaining any money for it!

Nevertheless I’m absolutely interested in any kind of responds - ideas, enhancements requests, criticism or bug reports. Call me, write me, mail me. I will hear you! YOU make Most alive and better for the future.

My address:

Uwe Roehm Woerthstr. 18 D - 94032 Passau Germany

EMail:

bugs@roemer.deg.sub.org

Fido:

2:2494/22.5

14 Current version

You can download the current version of Most from ftp.uni-passau.de (132.231.1.10) in the directory pub/amiga/local/UweR. You will also find other interesting software in pub/amiga/local…

I try to upload new releases of Most on AmiNet and on our local BBS "Die Box Passau" (2:2494/22.0) as soon as possible.

15 Greetings

Till now I worked about two years on Most. And still there’re old (and new bugs) to be found. If there weren’t the response from you, the user, I couldn’t make Most to what it is now. There’ve been many enhancements I included in new releases and a lot of bugs I removed. I can’t mention everyone who wrote me but want to list the most important ones:

Karl "Charly" Skibinski, David Göhler, Uli Zukowski, Stefan Sticht, Matthias "Tron" Scheler, Jan Holler, Dirk Federlein, Markus Uhlendahl

If you want to send me new translations for Most, please do this via snail mail or at least use uuencode sending an e-mail! I received some translations which I couldn’t use or unpack. Sorry! The current translations are done by:

− Paul van der Heu (holländisch) − Frank Geider (french) − Francesco Dipietromaria (italian) − Magnus Holmgren (swedish) − Antonio Joaquin Gomez Gonzalez (spanish)

A special thanks to Urban for his sugesstions and XPK. And a great thank you to everyone who wrote me bug reports, suggestions and of course sent me gifts!!!!

Index 33

Index

A

AINSI Sequences. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 AmigaGuide . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 AppIcon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 AppWindow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 ARexx. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 ARexx Port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 Author . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

B

Background. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

C

Clipboard . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 Commandline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Configuration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Copyright. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Criticism . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 Current version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

D

Default values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

E

Edit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 encrypted files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 Environments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Extras . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

G

Gadgets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 Giftware . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Greetings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32

H

Help . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

I

Ideas . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 Includefiles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 Info requester. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

K

Keyboard commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

M

Macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Memory managment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Menu Search. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Mouse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 Move . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

P

packed files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 Paging. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Parameter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Pipe support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 Processes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Project . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

R

Refreshing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

S

Scripts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 Search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 search requester . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 Searching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Settings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Shell start . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Startup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Suggestions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

T

Text managment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

Index 34

Text marking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 Thanks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 Title. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 ToolTypes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

V

Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

W

Window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 Workbench start . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

X

XPK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

Table of Contents

1 Copyright . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2 Background . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3 Configuration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

3.1 FILES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3.2 PORTNAME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3.3 PUBSCREEN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3.4 STARTUP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.5 WINDOWX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.6 WINDOWY. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.7 WINDOWW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.8 WINDOWH. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.9 SCREENMODE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.10 FONT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.11 FONTHEIGHT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.12 APPICON . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.13 ICONX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 3.14 ICONY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 3.15 BUFLEN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 3.16 PIPELEN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 3.17 QUOTES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 3.18 DIRS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 3.19 ENDS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 3.20 STACKSIZE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

4 Keyboard commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

4.1 Scrolling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 4.2 Searching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 4.3 Open file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 4.4 Info requester . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 4.5 Help . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 4.6 Edit file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 4.7 Quit file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

5 Menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

5.1 Project . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 5.2 Move . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 5.3 Menu Search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 5.4 Extras . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 5.5 Settings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

6 The main window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

7 The search requester. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

8 The AppIcon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

9 XPK support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

10 Pipe support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

11 ARexx Port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

11.1 GetFileList . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 11.2 GetScreenName . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 11.3 Quit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 11.4 Request. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 11.5 Set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 11.6 OpenNew . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 11.7 Open . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 11.8 Close . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 11.9 ToFront. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 11.10 GetWindowDims. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 11.11 Edit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 11.12 Info. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 11.13 FirstPage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 11.14 LastPage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 11.15 NextPage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 11.16 PrevPage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 11.17 NextLine . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 11.18 PrevLine. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 11.19 NextColumn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 11.20 PrevColumn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 11.21 NextTab . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 11.22 PrevTab . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 11.23 Goto. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 11.24 Search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

12 Known bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

13 Ideas - Bug reports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

14 Current version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

15 Greetings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32

Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33